Decimal equivalent of the binary number
101001.1011 is?
A 41.7875.
B 415875.
c 41.6875.
D 41.1875.
1
Expert's answer
2020-12-24T14:15:56-0500

do calculation from right to left leaving the decimal part first.

1*2^0 =1

0*2^1=0

0*2^2=0

1*2^3=8

0*2^4=0

1*2^5=32

total=1+8+32=41

now the calculation of the decimal part done from left to right

1*2^-1=0.5

0*2^-2=0

1*2^-3=0.125

1*2^-4=0.0625

total=0.6875

now combine the none decimal and the decimal part


=41.6875

the answer is c